The excluded volume for O2 gas is 0.0318 L per mole which is also four times the actual volume of the molecules. What is radius of O2 molecule?

Grade:

1 Answers

Vikas TU
5 years ago
Given, V’ –  V = 0.0318 L/mol
(R’^3 – R^3) = 3*0.0318/4*3.14............(1)
 
Also given,
V’ – V = 4V
or
V’ = 5V
R’^3 = 5R^3..............(2)
Putting value from eqn. (1) in (2) we get,
R^3 = 0.018988525
R = 0.2667 m.